### v2

# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

Abstract
In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs. TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each major task further contains three subtasks, while each subtask comprises 5,000 test samples. Given the inherent complexity of open molecule generation evaluation, we also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules. Our comprehensive benchmarking of 25 LLMs reveals the current limitations as well as potential areas for improvement in text-guided molecule discovery. Furthermore, we propose OpenMolIns, a specialized instruction tuning dataset established for solving challenges raised by TOMG-Bench. Fine-tuned on OpenMolIns, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench. Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Normally, molecule discovery is a trial-and-error process (Ekins, 2024), which requires repetitive experimentation and data analysis (Mattern & Grosser, 2023). Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market (Lee et al., 2018).
With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) (Wang et al., 2023), there has been a significant step forward. As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions. With the assistance of GNNs, researchers could analyse the properties of molecules (Cai et al., 2022) and generate new molecule candidates (Jin et al., 2018). However, challenges still exist. GNN-based methods struggle to generalize to different tasks (Chen et al., 2024), necessitating costly data collection and preparation for different downstream tasks. Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design (Li et al., 2024c).
In contrast, Large Language Models (LLMs) have shown their great generalization capability (OpenAI, 2023) and could be easily adapted to different research fields. For instance, Cascella et al. (2023) utilize ChatGPT for supporting clinical practice and Zhen et al. (2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
Figure 1: Comparison of Text-based Targeted Molecule Generation (a) v.s. Text-based Open Molecule Generation (b).
As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound (Weininger, 1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages. With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain (Li et al., 2024b). The generalization capability makes LLMs a viable option in molecule discovery. Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists (Zhang et al., 2024; Li et al., 2024d). They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging (Li et al., 2024a). A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation. Although the molecule-caption translation datasets including ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023) and L+M-24 (Edwards et al., 2024) are crucial for bridging the gap between the molecular and textual domains, they still face several limitations that need to be addressed:
On the one hand, there is a concern about the generalization of models trained on the molecule-caption translation task (Li et al., 2024b). In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task. In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples (Li et al., 2024b). This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well. On the other hand, a critical issue is the inability to propose new molecule structures. The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard. Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench). TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations. As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exactly matched molecule. Instead, we design automatic evaluation processes with chemical toolboxes like RDKit (Landrum, 2013) to test whether the generation meets the requirements. In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them. Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive assessment of whether LLMs truly grasp the molecular space. Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules. To solve the challenges raised by TOMG-Bench, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules with meticulously designed rules from an existing molecule database. OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and xlarge) to study the data scaling law of LLMs in the text-based open molecule generation task.
Our contributions are primarily threefold:
  • 1.
    TOMG-Bench introduces the task of Text-based Open Molecule Generation, which consists of three major tasks and their nine subtasks in total to synthetically benchmark the performance of LLMs in this task, demonstrating the limitations of the targeted molecule generation datasets like ChEBI-20.
  • 2.
    We benchmarked a suite of 25 LLMs, providing a brand-new perspective for the evaluation of LLMs.
  • 3.
    We introduce OpenMolIns, an instruction tuning dataset, enabling Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
2 Related Work
2.1 Development of AIs in Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022). Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules. Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods (Wigh et al., 2022; Wu et al., 2018; Zhou et al., 2023). For instance, MolReGPT  (Li et al., 2024b) leverages LLMs like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks. Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Networks (RNNs) (Grisoni et al., 2020; Popova et al., 2019), Graph Neural Networks (GNNs) (Wang et al., 2023; Sun et al., 2022), and Transformer-based networks (Xia et al., 2023; Balaji et al., 2023; Edwards et al., 2022).
2.2 Text-based Molecule Generation
Text-based Molecule Generation (Edwards et al., 2021) has recently emerged as a transformative task to molecule discovery. Previously, this task centres on generating targeted molecules based on natural language descriptions, requiring the creation of paired datasets of molecules and their corresponding textual representations. Early approaches leveraged transformer-based models like MolT5 (Edwards et al., 2022), employing self-supervised learning on large datasets to generate high-quality SMILES strings from textual inputs. Subsequent advancements, such as KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023), integrated molecular graphs or biochemical text to improve both understanding and generation capabilities. 3D-MoLM (Li et al., 2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations. Meanwhile, the application of LLMs like MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) as in-context learners has also shown significant promise. These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context. Most recently, MolReFlect (Li et al., 2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
3 TOMG-Bench
In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space. Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks. Additionally, we created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
Figure 2: Task illustrations of TOMG-Bench. Both molecules to the right of the arrow are acceptable solutions.
3.1 Dataset Categorization
The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want. As shown in Figure 2, following the difficulty of the tasks, we demonstrate the three basic tasks as well as their corresponding subtasks:
MolEdit emerges as the most straightforward task among the three domains, as an existing molecule is already provided, and LLMs are only required to make modifications to it, which tests the molecular structure knowledge of LLMs. In this case, we have crafted three subtasks for MolEdit: AddComponent, DelComponent, and SubComponent. In AddComponent, LLMs are instructed to add a specific functional group to the given molecule, and DelComponent challenges LLMs to remove a specified functional group from the provided molecule, while SubComponent is a hybrid of the previous two subtasks, requiring LLMs to first remove a designated functional group and then introduce a new one as specified to the molecule.
MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target. To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery and can be effectively validated: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics). These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
MolCustom is the most challenging task, where we have established three subtasks: AtomNum, BondNum, and FunctionalGroup. For AtomNum, LLMs are tasked with generating molecules that adhere to a specified count and type of atoms. BondNum involves the creation of molecules with a defined number and type of bonds. In FunctionalGroup, LLMs must generate a molecule that includes functional groups as specified. These subtasks may appear straightforward, yet they are deceptively challenging. They demand that LLMs have a sophisticated understanding of molecular syntax to precisely generate molecules that meet the complex criteria set forth.
Compared to previous text-based target molecule generation datasets, the task design in TOMG-Bench is rather naive, but focuses more on the details of the molecular substructure, which puts even more rigorous requirements on both the molecular understanding and generation capabilities of LLMs. Furthermore, the accurate understanding on molecule substructures and editing operations also ensures the explainability and trustworthiness of generated molecules.
3.2 Dataset Construction
Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations. For instance, ChEBI-20  (Edwards et al., 2022), a dataset for the molecule-caption translation task, contains only 26,407 samples for training. In contrast, datasets for the text-to-image generation task in the computer vision field like MS COCO  (Chen et al., 2015) have over 1,500,000 annotated captions on more than 330,000 images. Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming. In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction. Instead, we focus on basic molecule structural properties and molecule operations that could be validated by chemical toolboxes, which makes it easy to increase the volume of data, thereby resolving data hunger in chemical corpora.
In TOMG-Bench, three major tasks consist of nine subtasks in total. Each subtask is allocated 5,000 test samples. Figure 4 illustrates the data construction process.
For MolEdit and MolOpt, we sample molecules from a specific molecule database. We mainly consider two molecule databases: Zinc-250K (Sterling & Irwin, 2015), and PubChem  (Kim et al., 2019). Zinc-250K has 250,000 molecules, while PubChem has about 10 million molecules. Therefore, the data size of Zinc-250K is more suitable for us to build the test sets, while PubChem could be further sampled to generate the instruction tuning dataset. After sampling, we utilize RDKit (Landrum, 2013), a molecular informatics toolbox, to collect basic molecule statistics, including the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
For the MolCustom task, we randomly generate 5,000 instructions as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups. Further details are provided in the Appendix D.
3.3 Metric Design
The evaluation of the TOMG-Bench is facilitated through a set of carefully designed metrics tailored to the specific tasks within the benchmark.
Task (Weighted) Success Rate Similarity Novelty Validity
MolEdit
MolCustom
MolOpt
Table 1: All used metrics are listed here. ’✓’ means that a metric is calculated on a task.
3.3.1 MolEdit & MolOpt Evaluation
For the MolEdit and MolOpt tasks, as shown in Table 1, the following metrics are adopted:
Success Rate: We have developed automated testing processes with RDKit, which compare the generated molecule with the original molecule and check if the number of the specified functional group is correct after modification (for MolEdit task) and if the specified chemical property is optimized as required (for MolOpt task). Details of the automated testing process are shown in Appendix E. The success rate reflects the percentage of molecules that have successfully passed the rigorous testing processes. However, as it is impossible to track every change between the generated molecules and the original molecule (See Appendix E for detailed explanation), successfully passing the test processes does not necessarily mean that the generated molecule is modified from the original molecule.
Similarity: In this case, we need to evaluate the similarity between the original molecule and the generated molecule. When removing a specified functional group, the modification should follow the minimal change principle. Therefore, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps, instead of transforming the molecule into an entirely new one or inadvertently removing additional components. For implementation, we calculate the Tanimoto Similarity between the generated molecule and the original molecule by transforming them into the Morgan Fingerprints (Butina, 1999), which can be described by the following equation:
$\delta(m^{g},m^{o})=\frac{\left|fp_{m^{g}}\cap fp_{m^{o}}\right|}{\left|fp_{m^{g}}\cup fp_{m^{o}}\right|},$ (1)
where $m^{g}$ denotes the generated molecule, $m^{o}$ is the original molecule, and $\delta(m^{g},m^{o})$ is the Tanimoto Similarity between the generated molecule and the original molecule, while $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Then, the similarity score for a subtask $t$ in MolEdit and MolOpt, denoted as $\delta_{t}$, is determined by taking the average of the similarity scores from all individual test samples.
Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax. We use RDKit to parse the SMILES strings. Only the molecules that successfully pass the parser are considered valid molecules. A higher validity means that the model is more familiar with the grammar of molecule textual representations.
3.3.2 MolCustom Evaluation
For the MolCustom task, the metrics adopted are also shown in Table 1. As LLMs are required to generate the molecule from scratch, we employ a different testing process for the evaluation of success rate and introduce the novelty score instead of similarity:
Success Rate: For different subtasks in MolCustom, we have designed different testing processes by the chemical toolbox. We calculate the number of the specified atoms, functional groups, or bonds for the three subtasks to check if the generation meets the requirements. Detailed testing process of MolCustom could also be found in Appendix E. A high success rate indicates that the LLM is effectively generating molecules according to the request.
Novelty: However, since there may be numerous valid solutions, fulfilling the basic requirements alone does not guarantee the quality of generation. To enable the exploration of new molecule structures, we employ novelty as an indication of the quality, which compares the similarity between the newly generated molecules and existing ones in the database. The novelty score is particularly important for discovering potential drug or material candidates, where innovation can lead to significant advancements. For implementation, we choose the Zinc250K as the database and calculate the average Tanimoto Similarity between the generated molecules and the existing molecules. The novelty $n$ for the generated molecule $m^{g}$ can be calculated as follow:
$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc250K}\delta(m^{g},m^{k})}{\left|Zinc250K\right|},$ (2)
Then, the novelty score forthe subtask $t$ in MolCustom, denoted as $n_{t}$, is determined by taking the average of the novelty scores from all individual test samples.
Validity: Similarly, validity is also an important metric for the MolCustom task.
3.3.3 Average Weighted Success Rate
To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce an average weighted success rate to rank the performance of LLMs. Considering that the success rate alone can not ensure the accurate generation for MolEdit and MolOpt tasks and the novelty score for MolCustom is also a critical metric for the quality of molecule generation, we adopt similarity and novelty scores as weights of the success rate:
$W\!S\!R_{t}=\begin{cases}n_{t}\times S\!R_{t},&t\in\{MolCustom\}\\ \delta_{t}\times S\!R_{t},&t\in\{MolEdit,MolOpt\}\end{cases},$ (3)
where $W\!S\!R_{t}$ denotes the weighted success rate for a subtask $t$, while $\delta_{t}$ is the similarity score for the MolEdit and MolOpt tasks, $n_{t}$ represents the novelty score for the MolCustom tasks, and $S\!R_{t}$ is the corresponding success rate of the subtask. Then, the average weighted success rate $\overline{W\!S\!R}$ could evaluate the synthetic performance of LLMs among all the nine subtasks:
$\overline{W\!S\!R}=\frac{1}{9}\sum_{t}W\!S\!R_{t},$ (4)
where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of performance that considers both the success rate and quality of the generation.
3.4 OpenMolIns: Instruction Tuning Dataset
In this section, we introduce OpenMolIns, a specialized dataset derived from the PubChem database to help LLMs get familiar with text-based open molecule generation via instruction tuning. This instruction tuning dataset is meticulously designed to ensure that the molecules it contains do not overlap with those in the Zinc-250K dataset to promote the generation of more novel molecular structures and avoid any potential data leakage that could compromise the integrity of the model performance.
We collect the instruction tuning dataset by constructing the samples of the nine subtasks in equal amounts. The RDKit toolbox is applied to construct the instruction tuning dataset. Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal. Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit. In the case of MolOpt, we also evaluate the direction of the desired property changes by calculating the properties to determine whether the operation improves or decreases the property value. Then, the original molecule, the optimized molecule, and the optimization direction are altogether wrapped by a pre-defined prompt.
In the MolCustom domain, constructing the instruction tuning samples is rather straightforward. We calculate the molecular statistics and encapsulate them within prompts. For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers. Then, we wrap these statistics with a random pre-defined prompt to construct the training sample. This method allows the generated molecules to better fit the real distribution of molecular space.
To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels: light, small, medium, large, and xlarge, as illustrated in Section 3.5. Each level represents a different quantity of data, shown in Table 2, allowing researchers to analyze how the amount of training data influences the model’s ability to generate and edit molecules effectively.
Dataset TOMG-Bench OpenMolIns
item each subtask total light small medium large xlarge
Data Size 5,000 45, 000 4,500 18,000 45,000 90,000 1,200,000
Table 2: Statisics of TOMG-Bench and OpenMolIns.
3.5 Statistics
In this section, we outline the data sizes of TOMG-Bench along with the OpenMolIns dataset, which are shown in Table 2. For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples.
OpenMolIns has five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the text-based open molecule generation task. The nine subtasks within the TOMG-Bench are uniformly distributed in OpenMolIns.
4 Experiments
4.1 Models
The models benchmarked are categorized into four groups: proprietary models, open-source general LLMs, open-source ChEBI-20 fine-tuned LLMs, and OpenMolIns fine-tuned LLMs.
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful language understanding and generation capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), and chatglm-9B (GLM et al., 2024). These LLMs are top-tier open source general LLMs within our computation budget.
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large (Edwards et al., 2022), and BioT5-base (Pei et al., 2023). These models have been the state-of-art models in the molecule-caption translation task.
OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M. We select Galatica-125M because it has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small size model can also help us study the data scaling law within reasonable budget. Besides, Llama3.2-1B-Instruct and Llama-3.1-8B-Instruct are selected due to their advanced capabilities and similar architectures.
Figure 3: The performance of models benchmarked in TOMG-Bench. In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models of unknown parameters are denoted as horizontal lines.
4.2 Findings
Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns. More precise details for each subtask are further illustrated in Appendix F. Based on the above results, we observe the following key findings:
Text-based open molecule generation is both promising and challenging for LLMs. As illustrated in Figure 3, we have calculated the average weighted success rate across all the nine subtasks. Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving an average weighted success rate of 35.92%. Gemini-1.5-pro follows closely with an average weighted success rate of 34.80%. Meanwhile, fine-tuned on OpenMolIns-large, Llama-3.1-8B-Instruct achieves an average weighted success rate of 27.22%, outperforming all the open-source general LLMs.
However, these results still underscore the considerable scope for improvement, even among the most advanced proprietary LLMs. What’s more, despite LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge. In MolCustom tasks, no LLM has managed to achieve a success rate exceeding 25% for a single subtask. Meanwhile, as shown in Figure 5, among the three subtasks of MolCustom, BondNum is the most difficult task, as it requires LLMs to understand the connections or edges in the molecule graph merely from the textual representation. These observations indicate that the generation of molecules from scratch demands a deep understanding of the molecular structures, an area where current models are still striving to make significant strides.
Most powerful open-source general LLMs can already outperform GPT-3.5-turbo. In TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive average weighted success rate of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%. Compared to the proprietary models, open-source general LLMs seem to have been lack of chemical corpora for the pre-training process. However, recent development of open-source general LLMs have rapidly bridged the gap. The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and showing the increasing diversity of pre-training corpora.
More powerful LLMs usually inherit a better performance in TOMG-Bench. Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench. For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
Similarly, within the Llama-3 series, we could also observe that larger and more powerful models tend to achieve superior results on the TOMG-Bench. These findings underscore a strong correlation between an LLM’s capabilities and its performance in text-based open molecule generation.
However, we encountered an unexpected anomaly with certain open-source LLMs. Notably, Qwen2-7B-Instruct, despite its impressive ability to solve mathematical problems and its size of 7 billion parameters, underperformed models with as few as 1 billion parameters. This result is particularly striking and suggests that the TOMG-Bench offers a unique and comprehensive evaluation that current races for LLMs may not have adequately addressed. This discovery also highlights the significance of the TOMG-Bench as a new benchmark for LLMs. It provides a broader and more diverse assessment that exposes potential blind spots in the development of LLMs.
ChEBI-20 is insufficient for LLMs to master molecular structures and editing operations. The ChEBI-20 dataset is designed to bridge the gap between molecules and textual descriptions. Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in TOMG-Bench. For instance, BioT5-base, which is claimed as the state-of-the-art model for text-based molecule generation on the ChEBI-20 dataset, only achieves an average weighted success rate of 4.21% on the TOMG-Bench. In the MolEdit and MolOpt tasks, although these models obtain a high success rate, the generations are actually not modified from the original molecules, leading to poor similarity scores. This also shows that the average weighted success rate provides a more accurate reflection of the actual performance. Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a success rate above 5% in a single subtask. This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
The data scaling law varies among different tasks. Across five distinct data scales of OpenMolIns and the experiments on Galactica-125M, we observe a general data scaling law: as the size of the corpus increases, the performance of LLMs also improves. As shown in Figure 3 and Table 12, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
However, an interesting observation shows that the data scaling law varies among different tasks. For instance, in MolCustom tasks, we witnessed a substantial average improvement of 286% when we scaling the data size from large to xlarge, which suggests that, for specific tasks, increasing data size has a beneficial effect on model performance.
Conversely, in the MolCustom tasks, the performance gains from data scaling were notably modest. More concerning, in the MolEdit tasks, we even observed a decline in performance with increased data size, which indicates that the model’s capacity is not solely data-driven but is also constrained by the number of parameters. It suggests that for certain subtasks within TOMG-Bench, simply increasing the volume of data may not lead to proportional improvements in performance, and other factors such as model complexity or optimization strategies may need to be considered.
5 Conclusion
In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of LLMs in the realm of open-domain molecule generation. Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain. Additionally, we propose an instruction tuning dataset, OpenMolIns, to fully explore the potential of LLMs on the TOMG-Bench, enabling Llama-3.1-8B-Instruct to outperform GPT-3.5-turbo. Our contributions not only lie in the development of a novel benchmark for LLMs in molecular discovery but also provide a diverse indication of the LLM evaluation.
Appendix A Comparison with Existing Text-based Molecule Benchmarks or Datasets
Benchmarks or Datasets Task Type # Training Data # Test Data Public Access
PCdes(Zeng et al., 2022) Targeted Generation 10,500 3000 $\times$
ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
Mol-Instructions (description guided molecule design) (Fang et al., 2024) Targeted Generation 297,319 1000
TOMG-Bench & OpenMolIns Open Generation 1,200,000 45,000
Table 3: Comparison with existing text-based molecule generation benchmarks or datasets.
Table 3 shows the comparison between TOMG-Bench & OpenMolIns and the existing text-based molecule generation benchmarks or datasets. TOMG-Bench introduces the brand-new open-ended molecule generation task for LLMs, which boosts the diversity of text-based molecule generation task. Compared to the existing benchmarks or datasets, OpenMolIns has the largest training set, which could better aligns the molecular space with the natural language space. Meanwhile, TOMG-Bench owns three major tasks and nine subtasks with 45,000 test samples in total, whose test data size is also the largest one among all the existing benchmarks and datasets, enabling more accurate evaluation and comprehensive assessment of LLMs’ performance in text-based molecule generation.
Appendix B Implementation Details
We implement various scripts to facilitate the testing of the aforementioned models. For proprietary models, we adopt the OpenAI API [^1] https://platform.openai.com/docs/ framework. For open-source general LLMs, we utilize both the VLLM [^2] https://docs.vllm.ai framework and the OpenAI framework. For the remaining LLMs, we adopt the Hugging Face transformers library [^3] https://huggingface.co/docs/ for inference.
Furthermore, BioT5 is designed to use SELFIES, an alternative string representation of molecules by using special tokens to ensure valid generations (Krenn et al., 2019), as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
Appendix C Hyper Parameters
In this section, we illustrate the detailed parameters adopted in this work, as shown in Table 4.
Item Value
Generation
temperature 0.75
top_p 0.85
num_beams 1
max_new_tokens 512
Instruction Tuning
batchsize 32
lr 3e-4
cutoff_len 1024
Lora Settings
r 64
$\alpha$ 128
dropout 0.1
Table 4: Hyper-parameters.
Appendix D Data Construction
In this section, we introduce the construction details of TOMG-Bench and OpenMolIns dataset, as well as the prompt templates. Figure 4 demonstrates the overall data construction process for TOMG-Bench as well as the OpenMolIns. Generally, we utilize Zinc250K to construct the test samples in TOMG-Bench, while adopting PubChem for the training samples in OpenMolIns. Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
Figure 4: Data construction workflow of TOMG-Bench & OpenMolIns.
D.1 MolEdit
For the MolEdit task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs. In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent. Prompt templates for MolEdit are shown in Table 5. However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule. In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
Prompt Templates for MolEdit
AddComponent
Please add a {} to the molecule {}.
Modify the molecule {} by adding a {}.
Add a {} to the molecule {}.
DelComponent
Please remove a {} from the molecule {}.
Modify the molecule {} by removing a {}.
Remove a {} from the molecule {}.
SubComponent
Please substitute a {} in the molecule {} by {}.
Modify the molecule {} by replacing a {} by {}.
Replace a {} in the molecule {} by {}.
Please replace a {} in the molecule {} with {}.
Modify the molecule {} by substituting a {} with {}.
Substitute a {} in the molecule {} with {}.
Table 5: Prompt Templates for MolEdit.
Table 6 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights. Given that certain functional groups are encountered less frequently in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen.
Functional Group benzene ring hydroxyl aldehyde carboxyl amide
Weights 15 15 5 5 10
Functional Group amine nitro halo nitrile thiol
Weights 5 5 5 1 1
Table 6: Functional Groups that are considered in AddComponent and DelComponent, as well as their weights to be selected in AddComponent.
For SubComponent, our focus is exclusively on end groups for simplification, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
D.2 MolOpt
MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, TOMG-Bench requires one single LLM to optimize molecules with different metrics and optimization directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are listed in Table 7.
Prompt Templates for MolOpt
LogP
Please optimize the molecule {} to have a lower/higher LogP value.
Modify the molecule {} to decrease/increase its LogP value.
Optimize the molecule {} to have a lower/higher LogP value.
Please modify the molecule {} to decrease/increase its LogP value.
Modify the molecule {} to have a lower/higher LogP value.
MR
Please optimize the molecule {} to have a lower/higher MR value.
Modify the molecule {} to decrease/increase its MR value.
Optimize the molecule {} to have a lower/higher MR value.
Please modify the molecule {} to decrease/increase its MR value.
Modify the molecule {} to have a lower/higher MR value.
QED
Please optimize the molecule {} to have a lower/higher QED value.
Modify the molecule {} to decrease/increase its QED value.
Optimize the molecule {} to have a lower/higher QED value.
Please modify the molecule {} to decrease/increase its QED value.
Modify the molecule {} to have a lower/higher QED value.
Table 7: Prompt Templates for MolOpt.
Below are the detailed explanations of the chemical properties adopted in the three subtasks of MolOpt:
LogP refers to the logarithm of the partition coefficient, which is a measure of a molecule’s hydrophilicity or lipophilicity. It is an important factor in determining a compound’s bioavailability and membrane permeability.
Molecular Refractivity (MR) is a measure of the molar refractive index, which provides insight into the molecular size and the degree of molecular branching. It is used to assess the overall shape and bulk of a molecule.
Quantitative Estimation of Drug-Likeness (QED) is a computational metric that evaluates the drug-likeness of a molecule based on a set of predefined rules. A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
D.3 MolCustom
To enable customized design of molecules, we consider three fundamental features for describing the molecule, including atoms, bonds, and functional groups. Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request. The prompt templates for MolCustom are shown in Table 8. Below, we present the construction details of the three subtasks for MolCustom:
Prompt Templates for MolCustom
AtomNum
Please generate a molecule with {} atom(s).
Please generate a molecule composed of {} atom(s).
Please generate a molecule consisting {} atom(s).
The molecule has {} atom(s).
The molecule is composed of {} atom(s).
The molecule consists of {} atom(s).
There is a molecule with {} atom(s).
There is a molecule composed of {} atom(s).
There is a molecule consisting of {} atom(s).
The molecule contains {} atom(s).
BondNum
Please generate a molecule with {} bond(s).
Please generate a molecule composed of {} bond(s).
Please generate a molecule consisting {} bond(s).
The molecule has {} bond(s).
The molecule is composed of {} bond(s).
The molecule consists of {} bond(s).
There is a molecule with {} bond(s).
There is a molecule composed of {} bond(s).
There is a molecule consisting of {} bond(s).
The molecule contains {} bond(s).
FunctionalGroup
Please generate a molecule with {} group(s).
Please generate a molecule composed of {} group(s).
Please generate a molecule consisting {} group(s).
The molecule has {} group(s).
The molecule is composed of {} group(s).
The molecule consists of {} group(s).
There is a molecule with {} group(s).
There is a molecule composed of {} group(s).
There is a molecule consisting of {} group(s).
The molecule contains {} group(s).
Table 8: Prompt Templates for MolCustom.
Atom carbon oxygen nitrogen sulfur fluorine chlorine bromine iodine phosphorus
Weights [Mandatory] 5 3 3 2 2 2 2 1
Atom boron silicon selenium tellurium arsenic antimony bismuth polonium
Weights 1 1 1 1 1 1 1 1
Table 9: Atoms that are considered in AtomNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal atoms in the molecule structures.
AtomNum. Table 9 shows the atoms we consider in AtomNum, as well as their weights to be selected. Notably, carbon, as the basic unit in organic chemicals, is a mandatory option. The number of carbon atoms ranges from 1 to 40, while the number of other selected atoms ranges from 1 to 5. This setting relieves the difficulty for generation, as LLMs could generate a carbon backbone first and attach the remaining atoms to the backbone one by one.
Bond single double triple rotatable aromatic
Weights 5 4 3 1 1
Table 10: Chemical bonds that are considered in BondNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal bonds in the molecule structures.
BondNum. Similarly, we select five different kinds of chemical bonds: single, double, triple, rotatable, and aromatic, as shown in Table 10. For the single bond, if selected, the number can vary from 1 to 50. For the aromatic bond, the number follows the rules of the formation of aromatic bonds, varying from 5 to 20. Moreover, the number of these remaining bonds, if selected, is specified from 1 to 5.
Functional Group benzene ring hydroxyl anhydride aldehyde ketone carboxyl ester amide amine nitro
Weights 15 15 2 5 5 10 5 5 5 2
Functional Group halo thioether nitrile thiol sulfide disulfide sulfoxide sulfone borane
Weights 2 1 1 1 1 1 1 1 1
Table 11: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal and edge functional groups in the molecule structures.
FunctionalGroup. Lastly, we also specify functional groups in the molecule structure. Table 11 shows the range of functional groups and their weights that are taken into consideration.
Notably, in MolCustom, if not required, LLMs can generate any number of these unspecified atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
Appendix E Testing Process
In this section, we present the testing process used to calculate the success rate of each subtask.
Algorithm 1 MolEdit Testing Process
0:  Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Group-to-add $\alpha$, Group-to-delete $\Delta$
0:  Pass or Not (Bool) if $t$ is AddComponent then    if $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false else if $t$ is DelComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 then        return true     else        return false else if $t$ is SubComponent then    if $|\Delta|$($m^{g}$) = $|\Delta|$($m^{o})$ 1 and $|\alpha|$($m^{g}$) = $|\alpha|$($m^{o})$ + 1 then        return true     else        return false
The testing process for MolEdit task is demonstrated in Algorithm 1. However, in this process, we can not guarantee that the modified molecule differs the original molecule only in the specified functional group. The reasons are to follow:
  • The complexity of molecular structures makes it difficult to use regular matching to monitor the exact number of functional groups present. This is because molecules can have a wide variety of bonding patterns and spatial arrangements that are not always amenable to simple, rule-based matching techniques.
  • Modifying functional groups in specific positions can indeed lead to unintended changes in the molecule’s overall structure. This is due to the interconnected nature of atoms within a molecule, where altering one part can have a ripple effect on the stability and geometry of the entire structure.
  • The tools and algorithms currently available may not have the necessary sophistication to predict and control all the possible outcomes of a functional group modification, particularly when dealing with complex molecules or reactions that are not fully understood.
Therefore, the similarity score needs to be considered to ensure the accuracy and quality of the generation.
Algorithm 2 MolOpt Testing Process
0:  : Generated Molecule $m^{g}$, Original Molecule $m^{o}$, Subtask $t$, Optimization Direction Requirement $R$
0:  : Pass or Not (Bool) if $t$ is LogP then    if $LogP$($m^{g}$) $>$ $LogP$($m^{o})$ and R is $\uparrow$ then        return true     else if $LogP$($m^{g}$) $<$ $LogP$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is MR then    if $M\!R$($m^{g}$) $>$ $M\!R$($m^{o})$ and R is $\uparrow$ then        return true     else if $M\!R$($m^{g}$) $<$ $M\!R$($m^{o})$ and R is $\downarrow$ then        return true     else        return false else if $t$ is QED then    if $Q\!E\!D$($m^{g}$) $>$ $Q\!E\!D$($m^{o})$ and R is $\uparrow$ then        return true     else if $Q\!E\!D$($m^{g}$) $<$ $Q\!E\!D$($m^{o})$ and R is $\downarrow$ then        return true     else        return false
For MolOpt, we adopt Algorithm 2 for the testing process. Similarly, merely passing the test does not necessarily mean that the generated molecule follows the principle of minimal change.
Algorithm 3 MolCustom Testing Process
0:  : Generated Molecule $m^{g}$, Subtask $t$, Atom List $A$, Bond List $B$, Functional Group List $G$, Requirements $R$
0:  : Pass or Not (Bool) flag = trueif $t$ is AtomNum then    for $atom$ in $A$ do         if $|atom|$($m^{g}$) $\neq$ $R$[$atom$then            flag = false else if $t$ is BondNum then    for $bond$ in $B$ do         if $|bond|$($m^{g}$) $\neq$ $R$[$bond$then            flag = false else if $t$ is FunctionalGroup then    for $group$ in $G$ do         if $|group|$($m^{g}$) $\neq$ $R$[$group$then            flag = false return flag
For MolCustom, the testing process is shown in Algorithm 3, which test whether the specified number of the atom, bond, or functional group satisfies the requirements.
Appendix F Detailed Results
In this section, we first show the leaderboard of TOMG-Bench in Table 12, where Claude-3.5 achieves first place with an average weighted success rate of 35.92%. Notably, via instruction tuning on our OpenMolIns dataset, Llama-3.1-8B achieves 6th place, which outperforms all the existing open-source general LLMs.
Model #Parameters (B) $\overline{S\!R}$ (%) $\overline{W\!S\!R}(\%)$
Claude-3.5 (Anthropic, 2024b) - 51.10 35.92
Gemini-1.5-pro (Deepmind, 2024) - 52.25 34.80
GPT-4-turbo (OpenAI, 2023) - 50.74 34.23
GPT-4o (OpenAI, 2023) - 49.08 32.29
Claude-3 (Anthropic, 2024a) - 46.14 30.47
Llama-3.1-8B (OpenMolIns-large) 8 43.1 27.22
Galactica-125M (OpenMolIns-xlarge) 0.125 44.48 25.73
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 70 38.54 23.93
Galactica-125M (OpenMolIns-large) 0.125 39.28 23.42
Galactica-125M (OpenMolIns-medium) 0.125 34.54 19.89
GPT-3.5-turbo (OpenAI, 2023) - 28.93 18.58
Galactica-125M (OpenMolIns-small) 0.125 24.17 15.18
Llama3.1-8B-Instruct (Dubey et al., 2024) 8 26.26 14.09
Llama3-8B-Instruct (Dubey et al., 2024) 8 26.40 13.75
chatglm-9B (GLM et al., 2024) 9 18.50 13.13(7)
Galactica-125M (OpenMolIns-light) 0.125 20.95 13.13(6)
Llama3.2-1B (OpenMolIns-large) 1 14.11 8.10
yi-1.5-9B (Young et al., 2024) 9 14.10 7.32
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7 11.17 4.81
BioT5-base (Pei et al., 2023) 0.25 24.19 4.21
MolT5-large (Edwards et al., 2022) 0.78 23.11 2.89
Llama-3.1-1B-Instruct (Dubey et al., 2024) 1 3.95 1.99
MolT5-base (Edwards et al., 2022) 0.25 11.11 1.30(0)
MolT5-small (Edwards et al., 2022) 0.08 11.55 1.29(9)
Qwen2-7B-Instruct (Yang et al., 2024) 7 0.18 0.15
Table 12: Leaderboard of TOMG-Benchmark.
Then, we compare several representative LLMs from different model categories via radar maps, which are shown in Figure 5. Obviously, it can be concluded that Claude-3.5 outperforms almost all the other LLMs in these subtasks, except for the DelComponent task and the BondNum task. In DelComponent task, GPT-4o demonstrates better performance than Claude-3.5, while Llama-3.1-8B (OpenMolIns-large) achieves the best performance in the BondNum task.
Meanwhile, these radar maps also imply the difficulties of the subtasks. For example, for majority of the LLMs, they are struggling with the BondNum task. Even the most advanced proprietary LLMs like Claude-3.5 failed to achieve an ideal performance.
Moreover, in the MolOpt task, QED, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization. In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
Figure 5: The radar maps compare the performance ($\overline{W\!S\!R}$) of some representative LLMs from different model categories in each subtask of TOMG-Bench.
Finally, we present the detailed experimental results of all the subtasks in Table 13, 14, and 15.
Models AddComponent DelComponent SubComponent
SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
GPT-4o (OpenAI, 2023) 0.6188 0.6782 0.4197 0.7412 0.7012 0.6038 0.4234 0.8474 0.7992 0.7225 0.5774 0.9368
GPT-4-turbo (OpenAI, 2023) 0.6990 0.6936 0.4848 0.7934 0.7244 0.5735 0.4154 0.9060 0.7778 0.7323 0.5696 0.9160
GPT-3.5-turbo (OpenAI, 2023) 0.5832 0.6545 0.3817 0.7980 0.3082 0.7797 0.2403 0.8468 0.2918 0.6333 0.1848 0.6822
Claude-3.5 (Anthropic, 2024b) 0.6832 0.7017 0.4794 0.4414 0.5414 0.6678 0.3615 0.7960 0.8104 0.7310 0.5924 0.9588
Claude-3 (Anthropic, 2024a) 0.6766 0.6840 0.4628 0.8180 0.5556 0.6408 0.3560 0.8984 0.6550 0.7159 0.4689 0.9184
Gemini-1.5-pro (Deepmind, 2024) 0.7058 0.6792 0.4794 0.8254 0.7590 0.5949 0.4515 0.9158 0.7148 0.7139 0.5103 0.8684
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5198 0.6801 0.3535 0.5922 0.6122 0.5637 0.3451 0.7182 0.5094 0.7170 0.3652 0.6822
Llama3-8B-Instruct (Dubey et al., 2024) 0.3914 0.6649 0.2602 0.5374 0.4348 0.5058 0.2199 0.5700 0.2602 0.6841 0.1780 0.4838
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.2992 0.6088 0.1822 0.4954 0.4336 0.5257 0.2279 0.5910 0.3401 0.6424 0.2185 0.5076
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.1868 0.6251 0.1168 0.3760 0.2018 0.3774 0.0762 0.3590 0.0602 0.6227 0.0375 0.3550
Qwen2-7B-Instruct (Yang et al., 2024) 0.0010 0.2527 0.0003 0.0036 0.0006 0.4024 0.0002 0.0012 0.0004 0.2895 0.0001 0.0068
Yi-1.5-9B (Young et al., 2024) 0.1742 0.4170 0.0726 0.4216 0.2858 0.5936 0.1697 0.4909 0.1370 0.4619 0.0633 0.4368
Chatglm-9B (GLM et al., 2024) 0.2932 0.7622 0.2235 0.5686 0.2956 0.7494 0.2215 0.6914 0.1498 0.7150 0.1071 0.5084
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0374 0.5343 0.0200 0.1982 0.0768 0.5750 0.0442 0.3028 0.0102 0.3671 0.0037 0.1468
MolT5-small (Edwards et al., 2022) 0.1220 0.1027 0.0125 0.4490 0.1598 0.1125 0.0180 0.4504 0.0708 0.1029 0.0073 0.4876
MolT5-base (Edwards et al., 2022) 0.1354 0.1066 0.0144 0.4686 0.1562 0.1144 0.0179 0.4472 0.0584 0.1028 0.0060 0.4426
MolT5-large (Edwards et al., 2022) 0.2834 0.1084 0.0307 0.9282 0.2228 0.1201 0.0268 0.9198 0.1692 0.0932 0.0158 0.9410
BioT5-base (Pei et al., 2023) 0.3462 0.1567 0.0542 1.0000 0.1668 0.1597 0.0266 1.0000 0.0684 0.1576 0.0108 0.9998
Llama-3.2-1B (OpenMolIns-large) 0.1756 0.5676 0.0997 0.3216 0.1816 0.4963 0.0901 0.2466 0.0844 0.5415 0.0457 0.2958
Llama-3.1-8B (OpenMolIns-large) 0.5822 0.6541 0.3808 0.6730 0.5104 0.5074 0.2590 0.6896 0.5440 0.6258 0.3404 0.8400
Galactica-125M (OpenMolIns-light) 0.3786 0.5958 0.2256 0.6842 0.2062 0.6521 0.1345 0.7048 0.3102 0.5879 0.1824 0.6674
Galactica-125M (OpenMolIns-small) 0.3472 0.6172 0.2143 0.5356 0.3258 0.6025 0.1963 0.5758 0.2692 0.6181 0.1664 0.5692
Galactica-125M (OpenMolIns-medium) 0.4736 0.5682 0.2691 0.7442 0.4886 0.5184 0.2533 0.7488 0.3282 0.5975 0.1961 0.6958
Galactica-125M (OpenMolIns-large) 0.5866 0.5876 0.3447 0.8228 0.6078 0.5577 0.3390 0.7934 0.3438 0.6491 0.2232 0.8438
Galactica-125M (OpenMolIns-xlarge) 0.5842 0.5859 0.3423 0.8438 0.6526 0.5084 0.3318 0.8286 0.1872 0.6024 0.1128 0.8538
Table 13: Detailed results on MolEdit. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
Models LogP MR QED
SR Similarity WSR Validity SR Similarity WSR Validity SR Similarity WSR Validity
GPT-4o (OpenAI, 2023) 0.7190 0.6586 0.4735 0.8796 0.6864 0.6420 0.4407 0.8352 0.3952 0.6180 0.2442 0.8570
GPT-4-turbo (OpenAI, 2023) 0.7662 0.6984 0.5351 0.9048 0.7388 0.6821 0.5039 0.8848 0.3946 0.6587 0.2599 0.9050
GPT-3.5-turbo (OpenAI, 2023) 0.4048 0.6327 0.2561 0.8540 0.4120 0.6263 0.2580 0.8486 0.3316 0.5635 0.1869 0.8354
Claude-3.5 (Anthropic, 2024b) 0.7970 0.7124 0.5678 0.9422 0.6962 0.7112 0.4951 0.9110 0.5361 0.7042 0.3775 0.8604
Claude-3 (Anthropic, 2024a) 0.7984 0.6067 0.4844 0.9096 0.6094 0.6398 0.3899 0.9062 0.4678 0.5855 0.2739 0.9044
Gemini-1.5-pro (Deepmind, 2024) 0.7712 0.7022 0.5415 0.9274 0.7876 0.6744 0.5312 0.8926 0.4704 0.6077 0.2859 0.9484
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.5984 0.6028 0.3607 0.6482 0.5684 0.6032 0.3429 0.6272 0.2774 0.4828 0.1339 0.6340
Llama3-8B-Instruct (Dubey et al., 2024) 0.4642 0.3658 0.1698 0.6086 0.4332 0.4793 0.2076 0.5704 0.2568 0.4547 0.1168 0.6112
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.3990 0.4235 0.1690 0.5122 0.4336 0.5257 0.2279 0.5910 0.2655 0.4499 0.1194 0.6158
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.2220 0.4501 0.0999 0.2802 0.1908 0.2578 0.0492 0.3795 0.1210 0.3244 0.0393 0.2532
Qwen2-7B-Instruct (Yang et al., 2024) 0.0000 0.2923 0.0000 0.0004 0.0002 0.4123 0.0001 0.0004 0.0000 0.0000 0.0000 0.0000
Yi-1.5-9B (Young et al., 2024) 0.2884 0.5461 0.1575 0.4927 0.2050 0.3724 0.0763 0.4126 0.1064 0.6596 0.0702 0.4526
Chatglm-9B (GLM et al., 2024) 0.3666 0.6902 0.2530 0.4736 0.3514 0.6820 0.2397 0.5000 0.1832 0.6506 0.1192 0.4342
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0644 0.5055 0.0326 0.1664 0.0822 0.4410 0.0363 0.1604 0.0714 0.4757 0.0340 0.1796
MolT5-small (Edwards et al., 2022) 0.2158 0.1052 0.0227 0.4302 0.2316 0.1011 0.0234 0.4420 0.2214 0.1031 0.0228 0.4326
MolT5-base (Edwards et al., 2022) 0.2074 0.1051 0.0218 0.4168 0.1856 0.1073 0.0199 0.3796 0.2358 0.1054 0.0249 0.4536
MolT5-large (Edwards et al., 2022) 0.4244 0.1015 0.0431 0.8156 0.4496 0.1072 0.0482 0.8678 0.4654 0.1190 0.0554 0.9214
BioT5-base (Pei et al., 2023) 0.5158 0.1526 0.0787 1.0000 0.5060 0.1597 0.0808 1.0000 0.5068 0.1580 0.0801 1.0000
Llama-3.2-1B (OpenMolIns-large) 0.2898 0.5951 0.1725 0.3850 0.2644 0.5956 0.1575 0.3678 0.1996 0.5849 0.1167 0.3490
Llama-3.1-8B (OpenMolIns-large) 0.8054 0.6678 0.5378 0.8720 0.7122 0.6548 0.4663 0.8514 0.5224 0.6398 0.3342 0.8802
Galactica-125M (OpenMolIns-light) 0.3202 0.6547 0.2096 0.6416 0.3508 0.6435 0.2257 0.6358 0.2690 0.6521 0.1754 0.6380
Galactica-125M (OpenMolIns-small) 0.4172 0.6420 0.2678 0.5568 0.3958 0.6452 0.2554 0.5338 0.2956 0.6385 0.1887 0.5376
Galactica-125M (OpenMolIns-medium) 0.5904 0.5812 0.3431 0.7890 0.5874 0.5873 0.3450 0.7384 0.4608 0.5859 0.2700 0.7768
Galactica-125M (OpenMolIns-large) 0.6454 0.5927 0.3825 0.8198 0.6388 0.5973 0.3816 0.8028 0.4950 0.5962 0.2951 0.8100
Galactica-125M (OpenMolIns-xlarge) 0.7362 0.5744 0.4229 0.8902 0.7124 0.5697 0.4059 0.8612 0.5786 0.5677 0.3285 0.8626
Table 14: Detailed results on MolOpt. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).
Models AtomNum BondNum FunctionalGroup
SR Novelty WSR Validity SR Novelty WSR Validity SR Novelty WSR Validity
GPT-4o (OpenAI, 2023) 0.1998 0.6703 0.1339 0.5852 0.0650 0.6336 0.0412 0.8564 0.2330 0.6513 0.1518 0.8590
GPT-4-turbo (OpenAI, 2023) 0.1702 0.6991 0.1190 0.4904 0.0774 0.6301 0.0488 0.9068 0.2180 0.6605 0.1440 0.8778
GPT-3.5-turbo (OpenAI, 2023) 0.1070 0.5054 0.0541 0.6947 0.0518 0.6871 0.0356 0.5522 0.1136 0.6585 0.0748 0.8686
Claude-3.5 (Anthropic, 2024b) 0.1928 0.6926 0.1335 0.6548 0.1058 0.6584 0.0697 0.8860 0.2364 0.6582 0.1556 0.8892
Claude-3 (Anthropic, 2024a) 0.1044 0.6833 0.0713 0.5910 0.1042 0.6598 0.0688 0.8696 0.1816 0.9158 0.1663 0.6644
Gemini-1.5-pro (Deepmind, 2024) 0.1742 0.6902 0.1202 0.6774 0.0708 0.6522 0.0462 0.8688 0.2486 0.6673 0.1659 0.9240
Llama3-70B-Instruct (Int4) (Dubey et al., 2024) 0.1404 0.6675 0.0937 0.5474 0.0670 0.6478 0.0434 0.7378 0.1752 0.6576 0.1152 0.7650
Llama3-8B-Instruct (Dubey et al., 2024) 0.0242 0.6649 0.0161 0.3812 0.0260 0.6303 0.0164 0.5700 0.0848 0.6167 0.0523 0.7216
Llama3.1-8B-Instruct (Dubey et al., 2024) 0.0228 0.7020 0.0160 0.3862 0.0395 0.6541 0.0258 0.6387 0.1300 0.6274 0.0816 0.6905
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 0.0078 0.6732 0.0053 0.2986 0.0102 0.6309 0.0064 0.4524 0.0048 0.6012 0.0029 0.4020
Qwen2-7B-Instruct (Yang et al., 2024) 0.0110 0.9061 0.0100 0.2622 0.0010 0.8645 0.0090 0.0796 0.0022 0.8601 0.0019 0.0622
Yi-1.5-9B (Young et al., 2024) 0.0392 0.6848 0.0268 0.6170 0.0208 0.6407 0.0133 0.7072 0.0126 0.6945 0.0088 0.6521
Chatglm-9B (GLM et al., 2024) 0.0002 0.7483 0.0001 0.2131 0.0254 0.7189 0.0183 0.4682 0.0000 0.6908 0.0000 0.5926
Llama-3.2-1B-Instruct (Dubey et al., 2024) 0.0040 0.6807 0.0027 0.1850 0.0080 0.7465 0.0060 0.2226 0.0008 0.7461 0.0006 0.2818
MolT5-small (Edwards et al., 2022) 0.0006 0.6586 0.0004 0.6610 0.0064 0.5980 0.0038 0.6202 0.0114 0.5287 0.0060 0.8354
MolT5-base (Edwards et al., 2022) 0.0008 0.6868 0.0005 0.7560 0.0070 0.6509 0.0046 0.8422 0.0130 0.5464 0.0071 0.8382
MolT5-large (Edwards et al., 2022) 0.0150 0.7103 0.0107 0.8412 0.0118 0.5611 0.0066 0.8916 0.0382 0.6088 0.0233 0.9406
BioT5-base (Pei et al., 2023) 0.0118 0.8353 0.0099 0.9950 0.0078 0.6667 0.0052 0.9992 0.0476 0.6792 0.0323 0.9998
Llama-3.2-1B (OpenMolIns-large) 0.0144 0.6490 0.0093 0.5616 0.0350 0.6150 0.0215 0.6186 0.0252 0.6373 0.0161 0.4412
Llama-3.1-8B (OpenMolIns-large) 0.0136 0.6634 0.0090 0.7582 0.0544 0.6614 0.0360 0.7456 0.1344 0.6396 0.0860 0.6435
Galactica-125M (OpenMolIns-light) 0.0044 0.6054 0.0026 0.7930 0.0216 0.5724 0.0124 0.7596 0.0244 0.5756 0.0140 0.8442
Galactica-125M (OpenMolIns-small) 0.0146 0.6568 0.0096 0.8424 0.0530 0.6365 0.0337 0.7926 0.0570 0.5954 0.0339 0.8874
Galactica-125M (OpenMolIns-medium) 0.0294 0.6553 0.0193 0.8698 0.0622 0.6473 0.0403 0.7474 0.0882 0.6091 0.0537 0.8932
Galactica-125M (OpenMolIns-large) 0.0464 0.6729 0.0312 0.9116 0.0716 0.6695 0.0479 0.7374 0.0996 0.6276 0.0625 0.8966
Galactica-125M (OpenMolIns-xlarge) 0.1862 0.6899 0.1285 0.9308 0.1656 0.6887 0.1140 0.7952 0.2006 0.6445 0.1293 0.9162
Table 15: Detailed results for subtasks on MolCustom. For each task, we highlight and underline the best success rate (SR), as well as the weighted success rate (WSR).